# Reflecting with Two Voices: A Co-Adaptive Dual-Strategy Framework for LLM-Based Agent Decision Making

### v1

@@ Abstract
[…]
@@ Introduction
Large language models (LLMs) have shown promise in interactive task planning through reasoning and reflection (Yao et al. 2023b; Shinn et al. 2023). Yet, achieving robust performance in long-horizon, partially observable environments remains challenging. Most existing approaches rely on external demonstrations, retrieved exemplars, or auxiliary modules—leading to brittleness, limited generalization, high computational costs, and poor adaptability to novel situations.
@@
We evaluate DuSAR on ALFWorld (Shridhar et al. 2021) and Mind2Web (Deng et al. 2023) using open-source LLMs (7B–70B). Without any in-context examples, DuSAR achieves state-of-the-art results: 37.1% success on ALFWorld (Llama3.1-70B), outperforming Synapse (13.0%) and TRAD (9.9%), and 4.02% on Mind2Web—more than doubling the best baseline. Ablations confirm the necessity of dual-strategy coordination. Notably, optional integration of expert demonstrations into the Holistic Strategy yields an additional 3–4 percentage point improvement, highlighting DuSAR’s flexibility and compatibility with external knowledge. Our contributions are threefold:
@@
Early LLM agents combined Chain-of-Thought prompting with self-reflection to interleave reasoning and action in text environments (Yao et al. 2023b; Shinn et al. 2023; Yao et al. 2023a). Recent work extends this to interactive tasks—e.g., web navigation and embodied control—via tool/API integration (Brian Ichter 2023; Schick, Schramowski, and Sch”utze 2024; Qin et al. 2024). However, these approaches often falter in long-horizon settings due to challenges in dynamic replanning and partial observability, as evidenced in ALFWorld and WebArena (Shridhar et al. 2021; Zhu, Hernandez et al. 2023). Hierarchical methods improve scalability through task decomposition (Prasad et al. 2024; Liu, Du et al. 2024), but typically depend on manual designs or extensive demonstrations, limiting generalization. Critically, many rely on closed-source models (e.g., GPT-4) and degrade severely on open-source LLMs—ReAct (Yao et al. 2023c), for instance, achieves near-zero success without structured planning, hindering reproducibility and deployment.
@@
Current LLM agents often rely on external support—such as in-context demonstrations (Zhou et al. 2024; Zheng et al. 2024) or separate planner/executor modules (Liu, Du et al. 2024)—which limits their generalization to novel tasks and robustness under distribution shifts (Yoo, Kim, and Woo 2025). Retrieval-augmented methods like Synapse and TRAD require large demonstration banks, resulting in verbose prompts and high token overhead that hinder scalability and real-time use. Even self-reflection approaches (Kamoi et al. 2024) remain dependent on extensive context, incurring significant prompt costs (Wang, Zhang, and Li 2025). These dependencies underscore a critical gap: the lack of lightweight, zero-shot frameworks that can operate efficiently within a single frozen LLM while maintaining robustness in partially observable environments.
@@
To improve planning robustness, structured approaches have combined global LLM reasoning with local algorithmic modules (Cui et al. 2024) or hierarchical pipelines (Liu et al. 2025), inspired by the principle of separating stable planning from adaptive execution. While these methods introduce valuable modularity, they typically require external components, explicit supervision, or additional training—limiting their generality and efficiency. Crucially, they externalize the duality of stable planning and adaptive execution rather than internalizing it within a single model. Drawing inspiration from human metacognitive processes—where strategic planning and tactical execution continuously inform each other through reflective assessment—we abstract the principle of co-adaptive dual-strategy reasoning into a purely prompt-based framework, internalized within a single frozen LLM. DuSAR realizes this by instantiating two co-adaptive reasoning strategies (Holistic and Local Strategies) within a frozen LLM, enabling dynamic strategy refinement through Strategy Fitness Scores without external modules or training.
@@
To assess whether an LLM agent can achieve robust, generalizable planning without external demonstrations, additional training, or proprietary APIs—and at lower computational cost—we compare DuSAR against two state-of-the-art demonstration-driven agents: Synapse (Zheng et al. 2024) and TRAD (Zhou et al. 2024). Synapse uses single-stage trajectory-level retrieval based on task similarity, while TRAD employs a two-stage approach: first retrieving task-level exemplars, then refining based on intermediate reasoning (“thought”) similarity. Both rely on large external demonstration banks and are typically evaluated with proprietary models like GPT-4. In contrast, DuSAR requires no demonstrations and is designed for open-source LLMs.
We also include ReAct (Yao et al. 2023c) as a baseline that uses fixed in-context examples (two per task type). However, ReAct achieves negligible performance on open-source models (at most 1.49% on ALFWorld with Llama3.1-70B, and 0.13% on Mind2Web), leading us to omit its ALFWorld results from Table 1 due to effectively zero success across most model scales.
We evaluate all methods on two challenging benchmarks. ALFWorld (Shridhar et al. 2021) is a text-based household simulation requiring long-horizon planning, object affordance reasoning, and exploration under partial observability. Following standard protocol (Shridhar et al. 2021), we test on 134 out-of-distribution tasks across six sub-task categories—Put, Examine, Clean, Heat, Cool, and PutTwo—and report overall success rate (All SR).
Mind2Web (Deng et al. 2023) requires agents to complete real-world web tasks (e.g., booking a hotel) by interacting with raw HTML DOMs. We evaluate under three cross-generalization settings: Cross-Task, Cross-Website, and Cross-Domain, along with their macro-average (All). Metrics include Element Accuracy (Ele. Acc), Step-level Success Rate (Step SR), and Task Success Rate (Task SR). For fair comparison with prior work (Zhou et al. 2024), we summarize observations into 5 salient web elements using the pre-trained ranker from Deng et al. (Deng et al. 2023).
@@
Synapse 40.4 23.4 0.40 33.2 18.8 0.00 35.2 20.8 0.70 36.3 21.0 0.37
TRAD 42.9 37.8 3.17 34.8 28.1 0.56 38.7 33.9 2.14 38.8 33.3 1.96
DuSAR (Ours) 54.6 45.5 5.36 44.6 32.3 3.11 45.9 36.8 3.59 48.4 38.2 4.02
@@
As shown in Tables 2 and 3, DuSAR consistently outperforms all baselines in task success rate (Task SR) and element accuracy. With Llama3.1-70B, it achieves 4.02% Task SR—surpassing TRAD (1.96%), Synapse (0.37%), and ReAct (0.13%). While some baselines match DuSAR’s step-level accuracy on small models (e.g., Synapse: 9.2% vs. 9.7% on Llama2-7B), they lag far behind on Task SR, the key metric for multi-step completion, underscoring DuSAR’s stronger cross-step coherence. This gap widens under cross-domain evaluation, where DuSAR attains 3.00% (8B), 3.62% (12B), and 3.59% (70B) Task SR—versus TRAD’s 0.00–2.14%—demonstrating robust generalization without domain-specific exemplars.
Cross-Domain Generalization Challenges for Retrieval-Based Agents
@@
By eliminating demonstration retrieval, DuSAR maintains a compact reasoning trace. Reduced prompt length directly translates to lower latency and higher throughput in deployment scenarios. As Table 3 shows, it uses only 335–564 tokens per step—versus 1.5k–3.7k for Synapse/TRAD and 1.4k–1.9k for ReAct. This reduces context pressure and avoids error propagation from irrelevant exemplars. Crucially, DuSAR achieves superior Task SR (e.g., 4.02% vs. 0.13% on Llama3.1-70B) at a fraction of ReAct’s token cost, offering an exceptional trade-off between efficiency and performance.
@@
On ALFWorld, both OH and OL improve with model scale (0.075→0.28 and 0.060→0.29 SR from 8B to 70B), indicating that larger models can partially compensate for the absence of one reasoning pathway. However, NA—which combines both strategies without dynamic alignment—peaks at only 0.19 SR, far below full DuSAR (0.371 SR), demonstrating that co-adaptive integration is more critical than mere component presence.
On Mind2Web, the roles diverge: the Local Strategy (OL) scales effectively (0.14→0.36 Step SR), while the Holistic Strategy alone (OH) shows diminishing returns with scale (0.12→0.10 Step SR), suggesting that abstract planning requires local grounding to remain effective in complex, non-stationary web interfaces. Although NA improves with capacity (0.12→0.34 Step SR), it still lags behind the default co-adaptive design (0.46 Step SR). Crucially, on Llama3.1-8B, all ablated variants achieve at most 0.14 Step SR—less than half of DuSAR’s 0.34—confirming that robust performance stems from co-adaptive reasoning, not model scale alone.
@@
Results (Fig. 4) show that the effectiveness of demonstrations depends critically on model scale and the granularity of integration. On ALFWorld, high-level guidance (HT) provides substantial gains for smaller models: Llama3.1-8B achieves 29.8% SR (+18.1% over default DuSAR), as global structure compensates for limited planning capacity. In contrast, low-level traces (LT/BT) offer limited benefit at this scale (11.9%/10.4% SR), likely because fine-grained action sequences are difficult to contextualize without sufficient reasoning bandwidth. With Llama3.1-70B, however, BT achieves 46.2% SR (+9.1%), outperforming even HT (44.0% SR), demonstrating that large models can effectively integrate multi-granular demonstrations when reasoning capacity is abundant.
On Mind2Web, high-level guidance (HT) yields the highest Step SR for 8B (40.6% vs. 37.6% for LT and 34.1% for default), providing a modest 6.5% gain. All variants converge near 47–48% Step SR at 70B, yet default DuSAR already achieves 45.5% Step SR—and 4.02% Task SR—without any demonstrations.
Structured Reasoning Reduces Demonstration Dependence
@@ Conclusion
[…]
@@ 6 Opened cab 2: cup open cab 2 50 Opened
ReAct (Yao et al. 2023c) ✓ (fixed) Sequential reasoning-acting loop Reactive self-correction Medium (1.4k–1.9k/step)
Synapse (Zheng et al. 2024) ✓ (trajectory) Single-stage trajectory retrieval Task-based exemplar selection Moderate (1.5k–2.1k/step)
TRAD (Zhou et al. 2024) ✓ (step-wise) Two-stage retrieval (task + thought) Thought-based dynamic refinement Low (3.2k–3.6k/step)
DuSAR (Ours) Co-adaptive dual-strategy Dynamic co-adaptation High (335–564/step)
@@
In contrast, ReAct’s traces lack structured strategy evolution: each step generates independent reasoning without global coordination, leading to inefficient exploration patterns. For instance, ReAct may repeatedly click similar elements without recognizing task completion milestones, resulting in lower task success rates (0.13% vs. DuSAR’s 4.02% on Mind2Web with Llama3.1-70B).
@@
DuSAR (Llama3.1-70B, 5.36% success rate):
@@
  • Success Factor: Structured decomposition ensures all required fields are systematically addressed. Score milestones (30, 50, 80, 100) track form completion progress, preventing skipped fields. Token cost: 335–564 per step, 3–4× lower than retrieval methods.
@@
  • Success Factor: Cross-website UI pattern recognition enables adaptation to different search interfaces. Holistic Strategy provides flexible decomposition that adapts to website-specific layouts without relying on exemplar UI structures. Token cost: 335–564 per step.
@@
DuSAR (Llama3.1-70B, 3.59% success rate):
@@
  • Success Factor: Domain knowledge adaptation enables handling of specialized terminology and interfaces. Holistic Strategy decomposes medical/pharmacy tasks into domain-appropriate steps, while Local Strategy adapts to specialized UI elements without relying on domain-matched exemplars. Token cost: 335–564 per step, enabling efficient cross-domain generalization.
@@
  • Demonstration Requirements: Unlike Synapse (single-stage trajectory-level retrieval) and TRAD (two-stage step-level retrieval with thought-based refinement), DuSAR operates demonstration-free, enabling zero-shot generalization. ReAct (Yao et al. 2023c) uses fixed hardcoded examples (2 per task type) rather than dynamic retrieval, but still relies on external demonstrations. ReAct’s simple reasoning-acting loop struggles on open-source LLMs even with fixed examples: 1.49% SR on ALFWorld (Llama3.1-70B) vs. DuSAR’s 37.1%, and 0.13% SR on Mind2Web (Llama3.1-70B) vs. DuSAR’s 4.02%.
@@
  • Token Efficiency: DuSAR achieves 3–9× token reduction (335–564 vs. 1.5k–3.7k per step for retrieval methods) by eliminating demonstration overhead. Synapse incurs moderate token costs (1.5k–2.1k per step) due to single-stage trajectory retrieval. TRAD incurs higher token costs (3.2k–3.6k per step) due to two-stage retrieval (task similarity + thought similarity) and optional multi-step context expansion. ReAct uses moderate tokens (1.4k–1.9k per step) due to fixed hardcoded examples, but with limited performance on open-source models. DuSAR’s efficiency advantage is particularly pronounced in long-horizon tasks where retrieval methods accumulate demonstration overhead across steps.

### v2

@@ Abstract
Large language model (LLM) agents often rely on external demonstrations or retrieval-augmented planning, leading to brittleness, poor generalization, and high computational overhead. Inspired by human problem-solving, we propose DuSAR (Dual-Strategy Agent with Reflecting)—a demonstration-free framework that enables a single frozen LLM to perform co-adaptive reasoning via two complementary strategies: a high-level holistic plan and a context-grounded local policy. These strategies interact through a lightweight reflection mechanism, where the agent continuously assesses progress via a Strategy Fitness Score and dynamically revises its global plan when stuck or refines it upon meaningful advancement, mimicking human metacognitive behavior. On both simulated household (ALFWorld) and real-world web (Mind2Web) environments, DuSAR achieves state-of-the-art performance using only open-source LLMs, substantially outperforming all prior methods without any demonstrations or fine-tuning. Remarkably, it also reduces per-step token consumption by a large margin while maintaining strong task success. Ablation studies confirm the necessity of dual-strategy coordination. Moreover, optional integration of expert demonstrations further boosts performance, highlighting DuSAR’s flexibility and compatibility with external knowledge.
@@ Introduction
Large language models (LLMs) have shown promise in interactive task planning through reasoning and reflection [Yao et al., 2023b, Shinn et al., 2023]. Yet, achieving robust performance in long-horizon, partially observable environments remains challenging. Most existing approaches rely on external demonstrations, retrieved exemplars, or auxiliary modules—leading to brittleness, limited generalization, high computational costs, and poor adaptability to novel situations.
@@
We evaluate DuSAR on ALFWorld [Shridhar et al., 2021] and Mind2Web [Deng et al., 2023] using open-source LLMs (7B–70B). Without any in-context examples, DuSAR achieves state-of-the-art results: 37.3% success on ALFWorld (Llama3.1-70B), outperforming Synapse (13.0%) and TRAD (9.9%), and 4.00% on Mind2Web—more than doubling the best baseline. Ablations confirm the necessity of dual-strategy coordination. Notably, optional integration of expert demonstrations into the Holistic Strategy yields an additional 3–4 percentage point improvement, highlighting DuSAR’s flexibility and compatibility with external knowledge. Our contributions are threefold:
@@
Early LLM agents combined Chain-of-Thought prompting with self-reflection to interleave reasoning and action in text environments [Yao et al., 2023b, Shinn et al., 2023, Yao et al., 2023a]. Recent work extends this to interactive tasks—e.g., web navigation and embodied control—via tool/API integration [Brian Ichter, 2023, Schick et al., 2023, Qin et al., 2024]. However, these approaches often falter in long-horizon settings due to challenges in dynamic replanning and partial observability, as evidenced in ALFWorld and WebArena [Shridhar et al., 2021, Zhou et al., 2024b]. Hierarchical methods improve scalability through task decomposition [Prasad et al., 2024, Petruzzellis et al., 2025], but typically depend on manual designs or extensive demonstrations, limiting generalization. Critically, many rely on closed-source models (e.g., GPT-4) and degrade severely on open-source LLMs—ReAct [Yao et al., 2023b], for instance, achieves near-zero success without structured planning, hindering reproducibility and deployment.
@@
Current LLM agents often rely on external support—such as in-context demonstrations [Zhou et al., 2024a, Zheng et al., 2024] or separate planner/executor modules [Zhai et al., 2025]—which limits their generalization to novel tasks and robustness under distribution shifts [Yoo et al., 2025]. Retrieval-augmented methods like Synapse and TRAD require large demonstration banks, resulting in verbose prompts and high token overhead that hinder scalability and real-time use. Even self-reflection approaches [Kamoi et al., 2024] remain dependent on extensive context, incurring significant prompt costs [Wang et al., 2025b]. These dependencies underscore a critical gap: the lack of lightweight, zero-shot frameworks that can operate efficiently within a single frozen LLM while maintaining robustness in partially observable environments.
@@
To improve planning robustness, structured approaches have combined global LLM reasoning with local algorithmic modules [Wang et al., 2025a] or hierarchical pipelines [Liu et al., 2025], inspired by the principle of separating stable planning from adaptive execution. While these methods introduce valuable modularity, they typically require external components, explicit supervision, or additional training—limiting their generality and efficiency. Crucially, they externalize the duality of stable planning and adaptive execution rather than internalizing it within a single model. Drawing inspiration from human metacognitive processes—where strategic planning and tactical execution continuously inform each other through reflective assessment—we abstract the principle of co-adaptive dual-strategy reasoning into a purely prompt-based framework, internalized within a single frozen LLM. DuSAR realizes this by instantiating two co-adaptive reasoning strategies (Holistic and Local Strategies) within a frozen LLM, enabling dynamic strategy refinement through Strategy Fitness Scores without external modules or training.
@@
To assess whether an LLM agent can achieve robust, generalizable planning without external demonstrations, additional training, or proprietary APIs—and at lower computational cost—we compare DuSAR against two state-of-the-art demonstration-driven agents: Synapse [Zheng et al., 2024] and TRAD [Zhou et al., 2024a]. Synapse uses single-stage trajectory-level retrieval based on task similarity, while TRAD employs a two-stage approach: first retrieving task-level exemplars, then refining based on intermediate reasoning (“thought”) similarity. Both rely on large external demonstration banks and are typically evaluated with proprietary models like GPT-4. In contrast, DuSAR requires no demonstrations and is designed for open-source LLMs.
We also include ReAct [Yao et al., 2023b] as a baseline that uses fixed in-context examples (two per task type). However, ReAct achieves negligible performance on open-source models (at most 1.49% on ALFWorld with Llama3.1-70B, and 0.13% on Mind2Web), leading us to omit its ALFWorld results from Table 1 due to effectively zero success across most model scales.
We evaluate all methods on two challenging benchmarks. ALFWorld [Shridhar et al., 2021] is a text-based household simulation requiring long-horizon planning, object affordance reasoning, and exploration under partial observability. Following standard protocol [Shridhar et al., 2021], we test on 134 out-of-distribution tasks across six sub-task categories—Put, Examine, Clean, Heat, Cool, and PutTwo—and report overall success rate (All SR).
Mind2Web [Deng et al., 2023] requires agents to complete real-world web tasks (e.g., booking a hotel) by interacting with raw HTML DOMs. We evaluate under three cross-generalization settings: Cross-Task, Cross-Website, and Cross-Domain, along with their macro-average (All). Metrics include Element Accuracy (Ele. Acc), Step-level Success Rate (Step SR), and Task Success Rate (Task SR). For fair comparison with prior work [Zhou et al., 2024a], we summarize observations into 5 salient web elements using the pre-trained ranker from Deng et al. [Deng et al., 2023].
@@
Synapse 35.0 22.0 1.19 29.4 19.0 0.00 29.5 17.3 0.44 31.3 19.4 0.54
TRAD 44.2 37.6 2.38 36.0 29.0 0.85 38.8 33.5 1.94 39.6 33.4 1.75
DuSAR (Ours) 54.9 45.1 5.28 44.6 32.3 3.11 45.8 36.8 3.61 48.4 38.1 4.00
SR denotes Task Success Rate (Task SR), measuring complete task completion, distinct from Step SR which measures per-step action correctness.
@@
[…]
Generalization Challenges for Retrieval-Based Agents
@@
[…]
@@
On ALFWorld (All SR), both OH and OL improve with model scale (0.075→0.28 for OH; 0.060→0.29 for OL from 8B to 70B), indicating that larger models can partially compensate for the absence of one reasoning pathway. However, NA, which combines both strategies without dynamic alignment, peaks at only 0.19 SR, far below full DuSAR (0.373 SR). This demonstrates that co-adaptive integration is more critical than mere component presence.
On Mind2Web (Step SR), the roles diverge: the Local Strategy (OL) scales effectively (0.14→0.36 Step SR), while the Holistic Strategy alone (OH) shows diminishing returns with scale (0.12→0.10 Step SR), suggesting that abstract planning requires local grounding to remain effective in complex, non-stationary web interfaces. Although NA improves with capacity (0.12→0.34 Step SR), it still lags behind the default co-adaptive design (0.451 Step SR). Crucially, on Llama3.1-8B, all ablated variants achieve at most 0.14 Step SR—less than half of DuSAR’s 0.343—confirming that robust performance stems from co-adaptive reasoning, not model scale alone.
@@
As shown in Figure 4, the effectiveness of demonstrations depends critically on model scale and the granularity of integration. On ALFWorld, high-level guidance (HT) provides substantial gains for smaller models: Llama3.1-8B achieves 29.8% SR (+18.1 percentage points over default DuSAR), as global structure compensates for limited planning capacity. In contrast, low-level traces (LT/BT) offer limited benefit at this scale (11.9%/10.4% SR), likely because fine-grained action sequences are difficult to contextualize without sufficient reasoning bandwidth. On ALFWorld with Llama3.1-70B, however, BT achieves 46.2% SR (+8.9 percentage points), outperforming even HT (44.0% SR). This demonstrates that large models can effectively integrate multi-granular demonstrations when reasoning capacity is abundant.
[…]
Reduced Dependence on Demonstrations
@@ Conclusion
We present DuSAR, a demonstration-free, training-free framework that enables robust long-horizon planning in frozen open-weight LLMs through co-adaptive dual-strategy reasoning, which combines holistic subgoal decomposition with local, observation-grounded execution. DuSAR significantly outperforms retrieval-augmented baselines across simulation and web environments, all without demonstrations or fine-tuning. Crucially, this approach empowers even small models to generalize across domains via zero-shot abstraction, revealing that effective planning stems from architectural inductive bias rather than model capacity alone.
@@ 6 Opened cab 2: cup open cab 2 50 Opened
ReAct [Yao et al., 2023b] ✓ (fixed) Sequential reasoning-acting loop Reactive self-correction Medium (1.4k–1.9k/step)
Synapse [Zheng et al., 2024] ✓ (trajectory) Single-stage trajectory retrieval Task-based exemplar selection Moderate (1.5k–2.1k/step)
TRAD [Zhou et al., 2024a] ✓ (step-wise) Two-stage retrieval (task + thought) Thought-based dynamic refinement Low (3.2k–3.6k/step)
DuSAR (Ours) Co-adaptive dual-strategy Dynamic co-adaptation High (333–440/step)
@@
In contrast, ReAct’s traces lack structured strategy evolution: each step generates independent reasoning without global coordination, leading to inefficient exploration patterns. For instance, ReAct may repeatedly click similar elements without recognizing task completion milestones, resulting in lower task success rates (0.13% vs. DuSAR’s 4.00% on Mind2Web with Llama3.1-70B).
@@
DuSAR (Llama3.1-70B, 5.28% success rate):
@@
  • Success Factor: Structured decomposition ensures all required fields are systematically addressed. Score milestones (30, 50, 80, 100) track form completion progress, preventing skipped fields. Token cost: 335–564 per step (To_Pro), 3–4× lower than retrieval methods.
@@
  • Success Factor: Cross-website UI pattern recognition enables adaptation to different search interfaces. Holistic Strategy provides flexible decomposition that adapts to website-specific layouts without relying on exemplar UI structures. Token cost: 335–564 per step (To_Pro).
@@
DuSAR (Llama3.1-70B, 3.61% success rate):
@@
  • Success Factor: Domain knowledge adaptation enables handling of specialized terminology and interfaces. Holistic Strategy decomposes medical/pharmacy tasks into domain-appropriate steps, while Local Strategy adapts to specialized UI elements without relying on domain-matched exemplars. Token cost: 335–564 per step (To_Pro), enabling efficient cross-domain generalization.
@@
  • Demonstration Requirements: Unlike Synapse (single-stage trajectory-level retrieval) and TRAD (two-stage step-level retrieval with thought-based refinement), DuSAR operates demonstration-free, enabling zero-shot generalization. ReAct [Yao et al., 2023b] uses fixed hardcoded examples (2 per task type) rather than dynamic retrieval, but still relies on external demonstrations. ReAct’s simple reasoning-acting loop struggles on open-source LLMs even with fixed examples: 1.49% SR on ALFWorld (Llama3.1-70B) vs. DuSAR’s 37.3%, and 0.13% SR on Mind2Web (Llama3.1-70B) vs. DuSAR’s 4.00%.
@@
  • Token Efficiency: DuSAR achieves 3–9× token reduction (335–564 vs. 1.5k–3.7k per step for retrieval methods, To_Pro) by eliminating demonstration overhead. Synapse incurs moderate token costs (1.5k–2.1k per step) due to single-stage trajectory retrieval. TRAD incurs higher token costs (3.2k–3.6k per step) due to two-stage retrieval (task similarity + thought similarity) and optional multi-step context expansion. ReAct uses moderate tokens (1.4k–2.2k per step) due to fixed hardcoded examples, but with limited performance on open-source models. DuSAR’s efficiency advantage is particularly pronounced in long-horizon tasks where retrieval methods accumulate demonstration overhead across steps.